# Optimal Compressed Sensing for Image Reconstruction with Diffusion Probabilistic Models

### v2

@@ Abstract
We examine the problem of selecting a small set of linear measurements for reconstructing high-dimensional signals. Well-established methods for optimizing such measurements include principal component analysis (PCA), independent component analysis (ICA) and compressed sensing (CS) based on random projections, all of which rely on axis- or subspace-aligned statistical characterization of the signal source. However, many naturally occurring signals, including photographic images, contain richer statistical structure. To exploit such structure, we introduce a general method for obtaining an optimized set of linear measurements for efficient image reconstruction, where the signal statistics are expressed by the prior implicit in a neural network trained to perform denoising (generally known as a “diffusion model”). We demonstrate that the optimal measurements derived for two natural image datasets differ from those of PCA, ICA, or CS, and result in substantially lower mean squared reconstruction error. Interestingly, the marginal distributions of the measurement values are asymmetrical (skewed), substantially more so than those of previous methods. We also find that optimizing with respect to perceptual loss, as quantified by structural similarity (SSIM), leads to measurements different from those obtained when optimizing for MSE. Our results highlight the importance of incorporating the specific statistical regularities of natural signals when designing effective linear measurements.
@@ 1 Introduction
Natural signals, such as images and videos, are inherently high-dimensional. However, in many cases, imaging systems — such as cameras and microscopes — can only capture partial, lower-dimensional linear measurements of the target signal. Maximizing the informativeness of these measurements is critical for many real-world applications (Pruessmann et al., 1999; Zhu et al., 2018; Nehme et al., 2020; Deb et al., 2022). This raises a fundamental question: What are the optimal set of low-dimensional linear measurements for a given class of signals?
@@
To address this question, we first need to consider the task of reconstructing the original signal from a set of linear measurements, referred to as a “linear inverse problem”. In the general case, where the number of available linear measurements is smaller than the dimensionality of the signals to be reconstructed, linear inverse problems are under-determined. Solving an under-determined problem requires additional constraints, which, either explicitly or implicitly, impose assumptions about the structure of the signal distribution. Indeed, one can take a Bayesian approach to linear inverse problems, where the signal structure is expressed as a prior probability distribution over the signals. The inverse estimate is then computed from the posterior probability conditioned on the measurements, typically by minimizing an expected loss.
@@
For example, a widely-used approach to linear inverse problems is to assume a Gaussian signal prior, and mean squared error loss. In this case, the optimal set of linear measurements are the most significant principal components (PCs), and the optimal reconstruction is, in turn, achieved through linear projection onto those PCs. Although this approach is foundational, the Gaussian prior describes only the second-order statistical regularities of signals, and falls short of capturing important higher-order dependencies in natural many natural signal ensembles, particularly images.
@@
More recently, an important breakthrough in developing better linear measurements emerged by assuming sparse, heavy-tailed priors over natural images. The seminal work by Donoho (2006) in compressed sensing (CS) proved that when the signal lies within a union of subspaces (a type of sparse prior), the optimal measurements are incoherent with the axes of the signal subspaces, and are well-approximated by a set of randomly chosen vectors. In this setting, an iterative non-linear reconstruction from low-dimensional linear measurements can achieve near-perfect recovery (Tropp, 2006). The sparse prior is well-enough suited to some signal classes, such as medical images, that its use leads to significant empirical improvements. This advance highlights the importance of the signal prior in the design of effective linear measurements, as well as for the reconstruction algorithm.
@@
The statistical characterization underlying PCA and CS assumes alignment of signal content along principal component axes or subspaces, but this assumption does not adequately capture the statistical structure of natural images (Portilla et al., 2003b; Ballé et al., 2016). In support of this, previous work (Weiss et al., 2007) has demonstrated that random measurements outperform PCs for idealized sparse signals, but not for natural images. Over the past decade, deep neural networks (DNNs) trained for image processing tasks have been able to exploit ever more complex image statistical structure. Utilizing the priors implicit in these networks has led to remarkably improved performance in solving linear inverse problems (Romano et al., 2017b; Bora et al., 2017; Kadkhodaie & Simoncelli, 2020; Song et al., 2021). These dramatic performance improvements suggest that the network-based priors can more faithfully capture the statistical structure of natural images.
@@
Here, armed with the priors implicit in DNNs trained for denoising, hereafter a "network prior", we re-visit the question of optimal linear measurement for natural images. Specifically, we develop a framework for optimizing a set of linear measurements to minimize the error obtained via nonlinear image reconstruction under a network prior. More specifically, our reconstruction algorithm uses a generative diffusion model based on a trained DNN image denoiser. This enables us to apply our method to natural images and to analyze the impact of natural image statistics on the optimal linear measurements. We demonstrate that these measurements (1) vary substantially with the training dataset (e.g., digit vs. face images); (2) vary with the choice of reconstruction loss (e.g., MSE vs. SSIM); (3) are distinct from those of PCA, ICA and CS; and (4) lead to substantial performance improvements over PCA and CS. This work provides yet another example of the impressive improvements that can be achieved by applying modern ML methods to fundamental problems in signal processing. Our findings also establish a critical baseline for evaluating the potential benefits of non-linear measurements and the impact of measurement noise on the reconstruction of natural signals.
@@ 2.2 Image prior embedded in a denoiser
Traditionally, image priors were constructed by using simple parametric forms (Geman & Geman, 1984; Lyu & Simoncelli, 2008; Zoran & Weiss, 2011). Improvements in these priors led to steady progress over several decades. Over the last decade, however, the emergence of deep learning has made it possible to learn sophisticated priors from data. In particular, score-based diffusion models have exhibited incredible success in approximating image priors. Score-based diffusion models are deep neural networks trained to remove Gaussian white noise by minimizing mean squared error between the clean and denoised images. The learned denoiser is applied partially and iteratively, starting from a sample of Gaussian noise to generate an image. The generated image is a sample from the image prior embedded in the denoiser. The connection between denoising function and prior is made explicit in Tweedie’s equation (Miyasawa, 1961; Robbins, 1992; Raphan & Simoncelli, 2011; Efron, 2011):
@@
where $g_{\sigma}(z)$ is the distribution of Gaussian noise with variance $\sigma^{2}$. This is equivalent to convolution of $p(x)$ with the Gaussian probability density function. That is, $p_{\sigma}(y)$ is a blurred version of $p(x)$ where the extent of blur depends on $\sigma$. The family of $p_{\sigma}(y)$ forms a scale-space representation of $p(x)$ and is akin to the temporal evolution of a diffusion process of $p(x)$. A learned denoiser trained over a wide range of $\sigma$ approximates this family of gradient of log densities and is then used in a coarse to fine gradient ascent algorithm to sample from $p(x)$ (see Algorithm 1).
@@ 2.3 Inverse problem as constrained sampling
Images obtained using this algorithm are high-probability samples that are consistent with the measurements from the prior embedded in the learned denoiser. Notice that these sampling-based solutions to the inverse problem are not unique, and do not typically minimize mean square error: the MMSE estimate is a convex combination of these samples, and thus will not generally lie on the manifold of natural images from which the conditional samples are drawn. Visually, however, individual samples will look sharper and of higher visual quality compared to the MMSE estimate Kadkhodaie & Simoncelli (2021). In this work, we define the MMSE estimate as:
$\bar{x}(m)=h(m;M_{k})=\mathbb{E}_{x|m}[x]$ (4)
[…]
@@ 2.4 Optimized linear measurement
We aim to numerically find the set of $k$ linear measurements which minimizes average error of reconstruction through conditional sampling of the posterior. We define a loss function to measure the performance of our approximate posterior mean estimate, $\bar{x}(m)=h(m;M_{k})$:
@@
In order to solve the optimization problem of Equation 6, we use stochastic gradient descent in the space of all orthonormal matrices. Concretely, we parameterize the set of all orthogonal matrices $Q\in\mathcal{R}^{d\times k}$ using the Householder product, which represents matrices as a sequence of elementary reflections as the following (Trefethen & Bau, 2022; Shepard et al., 2015):
@@ 3.1 Two-dimensional example
Figure 1: Reconstruction of 2-D signals from a 1-D linear measurement. Top row: Gaussian distribution. Middle row: K-Sparse distribution. Bottom row: A data distribution confined to a curved one-dimensional manifold. A) Samples from the target distribution (blue points). Measurements are obtained via dot product between samples and a unit vector (black vector) parameterized by angle $\theta$. B) Mean squared error (MSE) of sample reconstruction as a function of measurement angle $\theta$, for two estimators: linear projection onto the measurement vector (blue), and reconstruction using the denoiser prior (orange). The dashed vertical line indicates the optimal solution, $\theta^{*}$, obtained with our optimization procedure (based on Equation 8). C) Test data drawn from the prior (blue), plotted in a new coordinate system $(x_{1}^{*},x_{2}^{*})$ rotated by $-\theta^{*}$ (i.e., such that the optimal measurement $m$ corresponds to the horizontal coordinate $x_{1}^{*}$). The bivariate estimates sampled from the denoiser prior conditioned on the optimal linear measurement are plotted in red. The MMSE solution depicted in the solid orange line is obtained by averaging over multiple conditional samples.
@@
We first consider a bivariate Gaussian distribution (Figure 1, top row). In this case, the first PC is the optimal measurement for both reconstruction methods, as expected for a Gaussian prior. The second row of Figure 1 depicts the result for a union-of-subspace sparse distribution. Here, the first PC is aligned with the horizontal axis. However, the optimal measurement vector for the denoiser prior is dramatically different. The reconstruction error at the optimal $\theta^{*}$ is near zero. This recapitulates the classical compressed sensing result (Donoho, 2006; Tropp, 2006) that for sparse signals, near-perfect recovery is achievable by making off-axis measurements.
The last row illustrates the scenario of primary interest in this paper: The data distribution lies in a low-dimensional but curved manifold. This type of higher-order structure cannot be adequately captured by either the Gaussian or sparse prior, but can be effectively described using the more powerful diffusion models, such as our denoiser prior (Supplemental Figure 1). Notably, in this case, we also identified an optimal measurement angle that outperforms the principal axis. Importantly, the optimal angle $\theta^{*}$ is correctly identified by our optimization method across all three cases, demonstrating the generality of our approach.
Finally, we highlight another aspect of the reconstruction demonstrated by the 2D examples (Figure 1C). Our methods generate estimates by sampling from the denoiser prior, conditioned on the linear measurements. While individual samples are consistent with the conditional prior, the nonlinear MMSE solutions are obtained by averaging multiple conditional samples. This averaging, however, can cause deviations of the MMSE from the prior. In the context of image reconstruction, the MMSE solution often exhibits lower visual quality (e.g., blurrier images) compared to individual conditional samples, as it does not lie close to the image prior manifold. This observation is particularly relevant in the results presented below and further motivates our approach of optimizing measurements with respect to a perceptual loss function later.
@@ 3.2 Optimized measurements for MNIST
Figure 2: Example image reconstructions using different measurement matrices. Columns show the PC projection, three conditional samples from the denoiser prior using the PCs, and three samples using the OLMs, respectively. Rows correspond to increasing number of measurements, $k$. The numbers indicate the PSNR value of the reconstructions, obtained by averaging over $16$ samples.
@@
We apply our method to find the optimized measurement for the MNIST dataset (Deng, 2012). To learn the prior distribution of these digit images, we train a neural network denoiser. For details of the architecture and training see Appendix E. We apply our method as described in Section 2.4 to obtain the optimized measurement matrix (OLM) for this dataset, for a range of $k$ values.
@@
For comparison, we choose two other types of linear measurements: the top $k$ PCs, and $k$ random vectors, which are the optimal measurements under Gaussian and sparse prior assumptions, respectively. The baseline error is simply that of linear projection onto the span of the PCs. Additionally, using the same $k$ PCs, we compute a non-linear reconstruction based on the denoiser prior. The top half of Figure 2 illustrates our results for a test digit image. First, we observe that combining the PCs with the denoiser prior significantly improves the linear inverse estimates which reflects the power of the denoiser prior. All conditional samples from the denoiser prior appear to be real MNIST images, which then gradually converge to the original as $k$ increases. Optimizing the measurement matrix offers additional improvements in the results, which reflects the importance of the measurements. These improvements are evident in both the identity of the digits and their more detailed appearances (see $k=4$ through $k=16$). Using random measurements (not shown) does not perform as well as either the optimized linear measurements or PCA, on average. See Supplemental Figure 2 for two more examples, including random measurement vectors.
@@
Figure 3: Performance of the optimized linear measurements, for A) MNIST and B) CelebA datasets. Left panel shows the peak signal-to-noise ratio (PSNR) as a function of the number of measurements $k$, for linear reconstruction from PCs (blue), denoiser prior reconstruction from PCs (orange), denoiser prior reconstruction from random measurements (green), and denoiser prior reconstruction from OLMs (red). Numbers indicate increase in PSNR from PC + denoiser prior compare to OLM + prior. The right panel shows a scatter plot comparing the PSNR values using the OLM (x-axis), and PCs (y-axis), over the images in the test set, for a value of $k$ around $10\%$ of the total number of pixels. Red points mark those shown in the examples of Figure 2 and Supplemental Figure 2 and Figure 3.
@@
We quantify our results in terms of peak signal-to-noise ratio (PSNR), as a function of $k$ in the left panel of Figure 3A. The OLMs result in superior reconstruction (red curve), particularly for the values of $k$ ranging from $1\%$ to $30\%$ of the total number of pixels. The peak performance gain from OLM over PC (both utilizing denoiser prior) reached 4.72 dB in PSNR at $k=256$. Additionally, we observe that PC measurements outperform random measurements using the denoiser prior (green), consistent with previous reports obtained with simpler priors (Weiss et al., 2007). This indicates that the union of subspace priors used in compressed sensing literature does not accurately describe the properties of natural images. Between the two PC reconstructions, the non-linear denoiser prior-based method (orange) shows a considerable improvement over linear reconstruction (blue), demonstrating the large advantage of using the more complex denoiser prior. See Supplemental Table 1 for comparison of our results to other methods in the literature.
@@
In addition to the average performance, we also compare the performance of these methods on individual images in the right panel of Figure 3A, in a scatter plot of the reconstruction errors from measurements using PCs versus that obtained using OLMs. In both cases, nonlinear reconstruction was performed using the denoiser prior; thus the plot highlights the differences due to the linear measurements. Importantly, we observe improvement in performance for almost all images in the test set for the OLM.

### v3

# Generalized Compressed Sensing for Image Reconstruction with Diffusion Probabilistic Models

@@ Abstract
We examine the problem of selecting a small set of linear measurements for reconstructing high-dimensional signals. Well-established methods for optimizing such measurements include principal component analysis (PCA), independent component analysis (ICA) and compressed sensing (CS) based on random projections, all of which rely on axis- or subspace-aligned statistical characterization of the signal source. However, many naturally occurring signals, including photographic images, contain richer statistical structure. To exploit such structure, we introduce a general method for obtaining an optimized set of linear measurements for efficient image reconstruction, where the signal statistics are expressed by the prior implicit in a neural network trained to perform denoising (known as a “diffusion model”). We demonstrate that the optimal measurements derived for two natural image datasets differ from those of PCA, ICA, or CS, and result in substantially lower mean squared reconstruction error. Interestingly, the marginal distributions of the measurement values are asymmetrical (skewed), substantially more so than those of previous methods. We also find that optimizing with respect to perceptual loss, as quantified by structural similarity (SSIM), leads to measurements different from those obtained when optimizing for MSE. Our results highlight the importance of incorporating the specific statistical regularities of natural signals when designing effective linear measurements.
@@ 1 Introduction
The natural world contains inherently high-dimensional spatial and temporal signals, but imaging systems - such as cameras and microscopes - typically capture only partial, lower-dimensional linear measurements of these signals. Maximizing the informativeness of these measurements is critical for many real-world applications (Pruessmann et al., 1999; Lustig et al., 2007; Zhu et al., 2018; Nehme et al., 2020; Deb et al., 2022). This raises a fundamental question: What is the optimal set of low-dimensional linear measurements for a given class of signals?
@@
To address this question, we first need to consider the task of reconstructing the original signal from a set of linear measurements, referred to as a “linear inverse problem”. In the general case, where the number of available measurements is smaller than the dimensionality of the signals to be reconstructed, the problem is underdetermined. Thus, obtaining a solution requires additional constraints that, either explicitly or implicitly, incorporate assumptions about the structure of the signal distribution. In the Bayesian formulation of the problem, the signal structure is expressed as a prior probability distribution over the signal space. This is combined with the measurement constraints (i.e., likelihood function) to obtain a posterior distribution, from which the reconstructed signal is computed, typically by minimizing an expected loss.
@@
In its most common form, one assumes a Gaussian signal prior, and mean squared error (MSE) loss. In this case, the optimal set of linear measurements are the most significant principal components (PCs), and the optimal reconstruction is the linear projection onto those PCs. Although this formulation is foundational, the Gaussian prior describes only the second-order statistical regularities of signals, and falls short of capturing important higher-order dependencies of many natural signal ensembles, particularly photographic images.
@@
More recently, an alternative framework for optimal linear measurements arose from the use of separable priors with sparse (heavy-tailed) marginals. The seminal work by Donoho (2006) in compressed sensing (CS) proved that when the signal lies within a union of subspaces (an idealized form of separable sparse prior), the optimal measurements should be incoherent with the axes of these subspaces, which can be implemented in practice using a set of randomly chosen vectors. In this setting, an iterative non-linear reconstruction can achieve near-optimal recovery (Tropp, 2006). The sparse prior provides a reasonable model for some signal classes, such as medical images, for which compressed sensing leads to significant empirical improvements over PCA methods. This advance highlights the importance of the signal prior in the design of effective linear measurements, as well as for the reconstruction algorithm.
@@
The statistical characterization underlying PCA and CS assumes alignment of signal content with a particular coordinate system, but this assumption does not adequately capture the statistical structure of natural images (Portilla et al., 2003b; Ballé et al., 2016). In support of this, previous work (Weiss et al., 2007) has demonstrated that although random measurements outperform PCs for idealized sparse signals, this does not hold for natural images. Over the past decade, deep neural networks (DNNs) trained for image processing tasks have been able to exploit ever more complex statistical structure, often described in terms of low-dimensional manifolds. Utilizing the priors implicit in these networks has led to substantial performance improvements in solving linear inverse problems (Romano et al., 2017b; Bora et al., 2017; Kadkhodaie & Simoncelli, 2020; Song et al., 2021). These dramatic performance improvements suggest that the DNN-based priors can more faithfully capture the statistical structure of natural images.
@@
Here, armed with the prior implicit in a DNN trained for denoising (hereafter a “denoiser prior”), we re-visit the question of optimal linear measurement for natural images. Specifically, we develop a framework for optimizing a set of linear measurements to minimize the error obtained via non-linear image reconstruction under the denoiser prior. More specifically, our reconstruction algorithm uses a generative diffusion model based on a trained DNN image denoiser. This enables us to apply our method to natural images and to analyze the impact of natural image statistics on the optimal linear measurements. We demonstrate that these measurements (1) vary substantially with the training dataset (e.g., digit vs. face images); (2) vary with the choice of reconstruction loss (e.g., MSE vs. SSIM); (3) are distinct from those of PCA, ICA and CS, with substantially more asymmetrical marginal distributions; and (4) lead to substantial performance improvements over PCA and CS. This work provides yet another example of the impressive improvements that can be achieved by applying modern ML methods to fundamental problems in signal processing. Our findings also establish a critical baseline for evaluating the potential benefits of non-linear measurements and the impact of measurement noise on the reconstruction of natural signals.
@@ 2.2 Image prior embedded in a denoiser
Traditionally, image priors were constructed by using simple parametric forms (Geman & Geman, 1984; Lyu & Simoncelli, 2008; Zoran & Weiss, 2011). Improvements in these priors led to steady progress over several decades. Over the last decade, however, the emergence of deep learning has made it possible to learn sophisticated priors from data. In particular, score-based diffusion models have exhibited incredible success in drawing samples from learned image priors. Score-based diffusion models are deep neural networks trained to remove Gaussian white noise by minimizing the mean squared error between the clean and denoised images. The learned denoiser is applied partially and iteratively, starting from a sample of Gaussian noise to generate an image. The generated image is a sample from the image prior embedded in the denoiser. The connection between denoising function and prior is made explicit in Tweedie’s equation (Miyasawa, 1961; Robbins, 1992; Raphan & Simoncelli, 2007; Vincent, 2011; Efron, 2011):
@@
where $g_{\sigma}(z)$ is the distribution of Gaussian noise with variance $\sigma^{2}$. This equation expresses the convolution of $p(x)$ with the Gaussian probability density function. That is, $p_{\sigma}(y)$ is a blurred version of $p(x)$ where the extent of blur depends on $\sigma$. The family of $p_{\sigma}(y)$ over all $\sigma>0$ forms a scale-space representation of $p(x)$ and is akin to the temporal evolution of a diffusion process of $p(x)$. A learned denoiser trained over a wide range of $\sigma$ approximates this family of gradients of log densities and can then be used in a coarse-to-fine gradient ascent algorithm to sample from $p(x)$ (see Algorithm 1).
@@ 2.3 Inverse problem as constrained sampling
Images obtained using this algorithm are high-probability samples that are consistent with the measurements from the prior embedded in the learned denoiser. Notice that these sampling-based solutions to the inverse problem are not unique, and do not typically minimize mean square error: the MMSE estimate is a convex combination of these samples, and thus will not generally lie on the manifold of natural images from which the conditional samples are drawn. The MMSE estimator is
$h(m;M_{k})\approx\mathbb{E}_{x|m}[x]$ (4)
which can be approximated by averaging over multiple conditional samples for a given measurement model $M_{k}$, see Section E.3 for more details. Visually, however, individual samples will look sharper and of higher visual quality compared to the MMSE estimate (Kadkhodaie & Simoncelli, 2021; Kawar et al., 2022).
@@ 2.4 Optimized linear measurement
We solve numerically for the set of $k$ linear measurements that minimize the average error of reconstruction through conditional sampling of the posterior. We define a loss function to measure the performance of our MMSE estimate, $h(m;M_{k})$:
@@
In order to solve the optimization problem of Equation 6, we use stochastic gradient descent in the space of orthonormal matrices. Specifically, we parameterize the set of matrices $Q\in\mathcal{R}^{d\times k}$ with orthonormal columns using the Householder product, which represents matrices as a sequence of elementary reflections as the following (Trefethen & Bau, 2022; Shepard et al., 2015):
@@ 3.1 Two-dimensional example
Figure 1: Reconstruction of 2-D signals from a 1-D linear measurement. Top row: Gaussian distribution. Middle row: K-Sparse distribution. Bottom row: A data distribution confined to a curved one-dimensional manifold. A) Samples from the target distribution (blue points). Measurements are obtained via dot product between samples and a unit vector (black vector) parameterized by angle $\theta$. B) Mean squared error (MSE) of sample reconstruction as a function of measurement angle $\theta$, for two estimators: linear projection onto the measurement vector (blue), and reconstruction using the denoiser prior (orange). The dashed vertical line indicates the optimal solution, $\theta^{*}$, obtained with our optimization procedure (based on Equation 8). C) Test data drawn from the prior (blue), plotted in a new coordinate system $(x_{1}^{*},x_{2}^{*})$ rotated by $-\theta^{*}$ (i.e., such that the optimal measurement $m$ corresponds to the horizontal coordinate $x_{1}^{*}$). Example bivariate estimates sampled from the denoiser prior conditioned on the optimal linear measurement are plotted in red. The MMSE solution (orange) is obtained by averaging over multiple conditional samples.
@@
We first consider a bivariate Gaussian distribution (Figure 1, top row). In this case, the first PC is the optimal measurement for both reconstruction methods, as expected for a Gaussian prior. The second row of Figure 1 depicts the result for a union-of-subspace sparse distribution. Here, the first PC is aligned with the horizontal axis. However, the optimal measurement vector for the denoiser prior is dramatically different. The reconstruction error at the optimal $\theta^{*}$ is near zero. This recapitulates the classical compressed sensing result (Donoho, 2006; Tropp, 2006) that for sparse signals, improved estimates can be achieved by making off-axis measurements.
The last row illustrates the scenario of primary interest in this paper: The data distribution lies in a low-dimensional but curved manifold. This type of higher-order structure cannot be adequately captured by either the Gaussian or sparse prior, but can be effectively described using the more powerful diffusion models, such as our denoiser prior (Supplemental Figure 2). Notably, in this case, we also identified an optimal measurement angle that outperforms the principal axis. Importantly, the optimal angle $\theta^{*}$ is correctly identified by our optimization method across all three cases, illustrating the generality of our approach.
Finally, we highlight another aspect of the reconstruction demonstrated by the 2D examples (Figure 1C). Our methods generate estimates by sampling from the denoiser prior, conditioned on the linear measurements. While individual samples are consistent with the conditional prior, the MMSE solutions are obtained by averaging multiple conditional samples. This averaging, however, can cause deviations of the MMSE solution from the prior. In the context of image reconstruction, the MMSE solution often exhibits lower visual quality (e.g., blurrier images) compared to individual conditional samples, as it does not lie close to the image prior manifold (Kadkhodaie & Simoncelli, 2021; Kawar et al., 2022). This observation is particularly relevant in the results presented below and further motivates our approach of optimizing measurements with respect to a perceptual loss function later.
@@ 3.2 Optimized measurements for MNIST
Figure 2: Example image reconstructions using different measurement matrices. Columns show the PC projection, three conditional samples from the denoiser prior using the PCs, and three samples using the OLMs, respectively. Rows correspond to increasing number of measurements, $k$. The numbers indicate the PSNR value of the reconstructions, obtained by averaging over $16$ samples. Red boxes highlight examples showing significant visual improvement of the OLM relative to the PC measurements. Top: denoiser trained on (and test example drawn from) the MNIST dataset. Bottom: denoiser trained on the CelebA dataset.
@@
We used our method to obtain optimized measurements for the MNIST dataset (Deng, 2012). We first trained a DNN denoiser on these digit images – for details of the architecture and training, see Appendix E. We apply our method as described in Section 2.4 to obtain the optimized measurement matrix (OLM) for this dataset, for a range of $k$ values.
@@
For comparison, we apply our reconstruction method to two other types of linear measurements: the top $k$ PCs, and $k$ random vectors, which are optimal measurements under Gaussian and sparse prior assumptions, respectively. We also compare to the optimal reconstruction for a Gaussian prior, which is simply linear projection onto the space spanned by the top $k$ PCs. The top half of Figure 2 illustrates our results for a test digit image. First, we observe that combining the PCs with the denoiser prior significantly improves the linear inverse estimates, which reflects the power of the denoiser prior. All conditional samples from the denoiser prior appear to be real MNIST images, which then gradually converge to the original image as $k$ increases. Use of the optimal measurement matrix (OLM) offers additional improvements, reconstructing the correct digit with a smaller $k$. These improvements are evident in both the identity of the digits and their more detailed appearance. Using random measurements (not shown) does not perform as well as either the optimized linear measurements or PCA. See Supplemental Figure 3 for two more examples, including random measurement vectors.
@@
Figure 3: Performance comparisons for A) MNIST and B) CelebA datasets. Left graph in each panel shows the peak signal-to-noise ratio (PSNR, in dB) as a function of the number of measurements $k$, for the denoiser prior reconstruction from OLMs (red), denoiser prior reconstruction from PCs (orange), linear (projection) reconstruction from PCs (blue), and denoiser prior reconstruction from random measurements (green). Inline numbers indicate the increase in PSNR from PC + denoiser prior to OLM + prior. The right graph in each panel plots the PSNR of the PC reconstruction against that of the OLM reconstruction over all images in the test set, for a value of $k$ approximately $10\%$ of the total number of pixels. Red points indicate the examples shown in Figure 2 and Supplemental Figure 3 and Figure 4. Also see Supplemental Table 1 for comparison of our results to other previous methods.
@@
We quantify our results in terms of peak signal-to-noise ratio (PSNR), as a function of $k$ in the left panel of Figure 3A. The OLMs result in superior reconstruction (red curve), particularly for the values of $k$ ranging from $1\%$ to $30\%$ of the total number of pixels. The peak performance gain of OLM over PC (both utilizing denoiser prior) reached 4.72 dB in PSNR at $k=256$. Between the two PC reconstructions, the nonlinear prior-based method (orange) shows a considerable improvement over linear reconstruction (blue), demonstrating the substantial advantage of using the more complex denoiser prior. See Supplemental Table 1 for a comparison of our results to other methods in the literature. Lastly, consistent with previous reports using simpler image priors (Weiss et al., 2007), we observe that PC measurements outperform random measurements when used with the denoiser prior (green). This raises the possibility that the union of subspace prior assumed in the compressed sensing literature (Donoho, 2006) does not fully capture the structure of these images, since random measurements are expected to perform well if the images indeed lie on a union of subspaces – a structure that we believe the denoiser prior is expressive enough to capture.
@@
In addition to the average performance, we also compare the performance of these methods on individual images in the right panel of Figure 3A, in a scatter plot of the reconstruction errors from measurements using PCs versus those obtained using OLMs. In both cases, nonlinear reconstruction was performed using the denoiser prior; thus the plot highlights the differences due to the linear measurements. Importantly, we observe improvement in performance for almost all images in the test set for the OLM.